OpenStudy (anonymous):

A moving van leaves a house traveling at an average rate of 40 mi/h. The family leaves the house 1/2hour later following the same route in a car. They travel at an average rate of 60 mi/h. a) Define the variable for the time traveled by the moving van. b) Write an expression for the time traveled by the car.

OpenStudy (anonymous):

r=d/t, rt=d, t=d/r. Those are your rate equations. See if you can use these to solve your problem.

OpenStudy (anonymous):

Yeah, but I don't know the distance or the time. Like for the moving van, my expression was x- 1/2 and for the car I put x. So, I multiplied them with the rate in a form of an equation..... 40(x- 1/2)= 60x .. However, when I solve it, I keep getting a negative answer, and I guess it's impossible to have a negative hour.

OpenStudy (anonymous):

For the van, time will equal d/40. for the car. The car has a time of d/60. If we want a time relative to when the the van started it will be t=d/60+.5 for all intervals over which t is larger than or equal to .5.